# David Roodman Open Philanthropy San Francisco, CA

### v2

@@ Abstract
Like Python and Java, which are integrated into Stata, Julia is a free programming language that runs on all major operating systems. The julia package links Stata to Julia as well. Users can transfer data between Stata and Julia at high speed, issue Julia commands from Stata to analyze and plot, and pass results back to Stata. Julia’s econometric software ecosystem is not as mature as Stata’s or R’s, or even Python’s. But Julia is an excellent environment for developing high-performance numerical applications, which can then be called from many platforms. The boottest program for wild bootstrap–based inference (Roodman et al., 2019) can call a Julia back end for a 33–50% speed-up, even as the R package fwildclusterboot (Fischer and Roodman, 2021) uses the same back end for inference after instrumental variables estimation. reghdfejl mimics reghdfe (Correia, 2016) in fitting linear models with high-dimensional fixed effects but calls an independently developed Julia package for tenfold acceleration on hard problems. reghdfejl also supports nonlinear models—preliminarily, as the Julia package for that purpose matures.
NOTE: This article is formatted for submission to the Stata Journal. It has not been accepted.
@@ 1 Introduction
In Stata, for example, a call to regress or gsem starts out in the ado programming language. The ado code translates your command line into a digital problem definition and passes that to routines written in C/C++ or Stata’s built-in Mata language. In modern Stata, an ado program can even call Python or Java. The called code will probably invoke the LAPACK library, which is written in Fortran, to invert matrices or solve linear systems.
@@
For creators of computationally intensive software, Julia has many strengths. More so than with Stata (but less so than with Python), if you need a tool to perform a fundamental task such as exploiting GPUs, there’s a good chance someone has written and posted it. Julia’s creators set out to solve the “two-language problem,” the inefficiency that arises when programmers have to split a project between high- and low-level languages as they trade off ease of software development and maintenance for speed of execution. (Python and C++ is a common combination, as is R and C++.) Julia resembles Python in syntax, but was built from the start for just-in-time compilation, meaning that it translates Julia functions to machine language when they are first called, and stores that code for reuse. Like LISP, a language almost as old as Fortran, Julia can treat code as data. This feature opens the door to meta-programs that can, for example, automatically optimize one’s code in certain ways by rewriting it before compilation, or generate new code to take derivatives of complex calculations already implemented.
@@
Julia has disadvantages. While the Julia ecosystem contains the foundation for statistical work—packages to manage data sets, fit standard models, and generate plots—many packages are immature or poorly documented. Standards for passing results are less well developed than in Stata. Much less statistical functionality is available in Julia than in Stata or R, or even Python.
@@
Since Julia is a great home for implementing and optimizing numerical software, but not (yet) a great home for the end user, a sweet spot for the new language is back end development. Methods of estimation and inference can be implemented in Julia, and called from Stata, R, and other platforms. The core work can implemented, optimized, and augmented in one place, rather than separately on each platform.
@@
The julia package links Stata to Julia. It follows in the footsteps of many projects to link statistical software to other languages. StataCorp introduced the Stata plugin interface (SPI) for C/C++ in Stata 8.1, in 2003. It added the built-in Mata language in Stata 9. Fiedler (2012) envisioned running Python inside of Stata. Following the suggestion of a StataCorp employee, Fiedler (2013) partly realized that vision by writing a plugin for Stata. Perhaps inspired by Fiedler, Stata 16 fully integrated Python. Java plugins became possible in Stata 13 and the Java prompt debuted in the Stata results window in version 17. Haghish (2019) introduced rcall, which allows one to issue R commands from Stata; it passes content between the two systems by saving files to disk. Outside of Stata, the JuliaCall package (Li, 2019) lets R users call Julia routines. An identically named package does the same for Python, and is paired with PyCall for Julia. [^1] See juliapy.github.io/PythonCall.jl/stable.
@@
The julia package was most inspired by Fiedler (2013). Its core too is a C++ plugin, which comes precompiled for Windows, Linux, macOS/Intel, and macOS/ARM. From the user’s point of view, the package has three components:
@@
The package is designed to hide the complexities of the Stata-Julia linkage, so that getting started requires just two quick steps, installing Julia and installing the julia package.
Two user-written Stata packages call Julia via julia. The boottest program for wild bootstrap–based inference (Roodman et al., 2019) now accepts a julia option, which causes it to use my Julia package WildBootTests.jl rather than the Mata back end. And the new reghdfejl mimics reghdfe and ivreghdfe, programs for fitting linear models with high-dimensional fixed effects (HDFE) (Correia, 2016), while calling FixedEffectModels.jl, whose lead developer is Matthieu Gomez. reghdfejl also has the ability, in beta, to fit nonlinear HDFE models, such as the logit and negative binomial, via Johannes Boehm’s GLFixedEffectModels.jl. In the last case, using Julia brings new functionality to Stata. In all cases, calling Julia cuts runtime on hard problems, sometimes by an order of magnitude.
Stata users can thus benefit from Julia and julia without using them directly. Still, this paper explains the new julia package. So it starts with a hands-on demonstration of running Julia from Stata, in section 2. Then section 3 demonstrates applications: it introduces reghdfejl and the Julia back end for boottest; and it demonstrates graphing via Julia. Section 4 gives tips and warns of some limitations. Section 5 more thoroughly describes the package’s Stata commands and Julia functions.
@@ 2.1 Installation and first examples
The preferred way to get Julia is to download the juliaup installation manager. You do that in Windows by installing “Julia” (not “juliaup”) from the Windows Store. [^3] The instructions might change. See github.com/JuliaLang/juliaup#installation for the latest. In Linux and macOS, you can instead start Stata and type
@@
The output may take several minutes to appear. On first use, jl installs and precompiles two Julia packages, DataFrames.jl and CategoricalArrays.jl, along with all the packages they depend on. DataFrames.jl (Bouchet-Valat and Kamiński, 2023) is the dominant Julia framework for data sets—i.e., rectangular grids whose columns have names and potentially different data types. CategoricalArrays.jl implements the Julia equivalent of Stata variables with data labels—ones coded, for example, so that 0 means “black” and 1 means “white.” Both packages are hosted on GitHub and are maintained by users, who accept improvements from others.
In general when using Julia, you will experience delays at four points: when packages are installed; when they are first used; when they are first used in a Julia session; and when particular features within them are first used, such as clustering standard errors. A thrust of recent development in the Julia langauge has been to move the behind-the-scenes processing from later to earlier stages in order to shorten the lags during regular use.
In the session shown below, I load Stata’s venerable “auto” data set, extract five variables, run a regression on them, and then copy the data to a Julia DataFrame named auto:
@@
jl save is not a Julia command. It is a subcommand of the Stata command jl, which is part of the julia package. If an object named auto already exists in the Julia environment, it is overwritten.
Next I type jl by itself, which changes the prompt in much the same way that typing mata or python or java does. Have started an interactive Julia session, I display the new DataFrame by typing its name:
@@
Along the top are the column names, and below them the types. Int16 means a 16-bit integer, equivalent to Stata int. Float32 corresponds to Stata float. The question marks at the ends of the type names indicate that the types have been modified to allow the special value missing. By default, jl save made all the columns accept missing. jl save also converted the foreign variable, which is coded as 0/1 in Stata and has value labels “Domestic” and “Foreign,” to a CategoricalVector. While the Julia version of the variable also has an internal, integer representation, the vector presents as a text variable. [^4] Julia categorical variables are not quite comparable to Stata variables with data labels. Code involving the Julia variable refers to display values, such as “Domestic,” not to numerical codes. The type label “CategoricalVector?” does not fit in the column head, so it is shortened to “Cat…?”.
Next I run the same regression in Julia, with the reg() function from the Julia package FixedEffectModels.jl:
  • jl> . using FixedEffectModels
@@
  • Stata has the # and ## interaction operators. The corresponding symbols in Julia are & and *. The first appears here.
  • All variable names in the command line are prefixed with “:”. That indicates that the names are symbols, not stand-alone Julia objects like the DataFrame auto.
  • Because foreign is a CategoricalVector, reg() automatically treats it as a factor variable, generating a pair of dummies to be interacted with mpg. There is no need for an “i.” prefix or the equivalent as in Stata. However, treating non-CategoricalVectors as factor variables requires a more awkward syntax. One must pass reg() an argument with a dictionary mapping the variables to coding types, such as with contrasts=Dict(:headroom=>DummyCoding()).
The command line contains deeper complexity. In Julia, as in R and Python, every regression command is a statement in a full-fledged programming language. This tends to complicate the syntax for even basic operations. It is hard to implement an R or Python line as clean as regress y x. In the Julia command above, term(:price), term(:headroom), …, etc., create instances of the object type Term to represent variables within a statistical model. The Term structure is defined in the StatsModels.jl package, which FixedEffectModels.jl loads and surfaces for the user. The content of each Term is just a symbol, such as :price. What give Term’s life are the operations that StatsModels.jl defines on them. The + operator is repurposed (“overloaded”) to form tuples of Term’s—pairs, triplets, and so on. The $\sim$ operator binds together two Term tuples to make a FormulaTerm, a structure whose distinctive feature is having designated left and right sides. The entire FormulaTerm is passed as the second argument to reg(). In this way, the primitives of a flexible programming language are composed to communicate a statistical model.
This standard for representing regression models is not intrinsic to Julia. It has emerged from the work of users. And it is transparent. In Stata, one could type jl: dump(term(:c) $\sim$ term(:a) + term(:b)) to reveal the contents of a FormulaTerm. The software that builds and applies these structures is on GitHub, all written in Julia. The StatsModels.jl formula language is also extensible [^5] See juliastats.org/StatsModels.jl/stable/internals/#extending.. As we will see, FixedEffectModels.jl defines a way to mark variables whose fixed effects are to be absorbed.
Yet a powerful feature of Julia makes it practical to ignore most of this complexity. The regression command above can be simplified with a macro. Where macros in Stata are strings, in Julia they are code for manipulating code—meta-programs. StatsModels.jl defines @formula, which lets one write models in a more R-like way. Thus:
@@
Before this new command is compiled and executed, @formula turns it into something closer to the previous one. [^6] Johannes Boehm’s experimental Douglass.jl package includes macros to translate Stata-like commands into Julia ones. An example: d"bysort :Species : egen :z = sum(:SepalLength) if :SepalWidth .> 3.0".
Notice one other change to the command line: it now starts with m= in order to store the fitted model in the Julia variable m. This lets me extract return values and pass them back to Stata:
@@ 2.2 Built for speed
To partially convey why Julia is a good environment for developing numerical software, I present another example. Suppose we have a data set with 10 million rows and 10 columns named x1-x10. Given a $10\times 10$ matrix ${\bf Q}$, we want to compute the norm of each row according to the quadratic form defined by ${\bf Q}$ and store the results in a new column. That is, for each row $x_{i}$, now viewed as a column vector, we want $x_{i}^{\prime}{\bf Q}x_{i}$. It is natural to implement this computation with triply nested for loops. That’s a good way to do it in Julia (inside Stata):
@@
We saw in the previous example that one can start an interactive Julia session inside Stata by typing jl (and end it with exit()). But the interactive mode is not accessible to do files. To exemplify Julia code in a do file, the above snippet is therefore written effectively as a single line. It is typographically split across multiple lines with Stata’s continuation token, ///. When putting many Julia commands on what is syntactically one line, most commands not immediately followed by end need to terminate with a semicolon.
The code for this new Julia function, XQX(), is mostly straightforward. It initializes the return value to a column of 0’s, performs the calculation through nested loops, and returns the result. One detail is unusual: another macro is invoked. @tturbo is part of Chris Elrod’s LoopVectorization.jl package. @tturbo analyzes the code that follows it and substantially—but invisibly—rewrites it for speed. It may unroll a loop to reduce the number of jumps back to the top. The macro, or the compiler, may spot the opportunity to move X[i,j] * out of the inner loop since it does not depend on the loop’s index, k. The macro may even reorder the nesting of the loops with an eye toward the fact that in Julia, matrices are stored column by column, and on-chip memory caching makes it is faster to access adjacent memory locations in sequence. The LoopVectorization package is called that because it also tailors code so that the resulting machine language exploits the ability of modern chips to vectorize operations within each core, in 256- or 512-bit registers—what is called single instruction, multiple data (SIMD) execution. And @tturbo will exploit multithreading, the subdivision of work across cores; that is what the first “t” in tturbo stands for.
Triply nested implementations in Mata and Python are much slower, I assume because the calculations in the innermost loop have to be interpreted a billion times. The best alternatives I have found in the two languages are essentially the same:
@@
Both delegate the computation to built-in matrix operations, which are fast because they are implemented in compiled C or Fortran. However, both contain a subtle inefficiency. They create the large, temporary matrices ${\bf XQ}$ and ${\bf XQ\otimes\bf X}$. Since in our example ${\bf X}$ is $10^{6}\times 10$, and since double-precision numbers take 8 bytes, the temporary matrices require some 800 megabytes each. Allocating and deallocating memory itself takes time. Expanding the memory footprint of an algorithm can also reduce speed by exceeding the capacity of a CPU’s memory caches, forcing the CPU to idle while data are written to and read from the computer’s main memory. The triply-looped Julia code demands essentially no temporary storage.
To create a test bed for these functions, I construct ${\bf X}$ as a Stata data set and ${\bf Q}$ as a Stata matrix, then copy them into Mata, Python, and Julia. The set rmsg on command generates timings:
@@
Under a Julia convention, the semicolons at the end of the jl lines suppress the printing of the results of the assignments. The Julia data copying command is run twice in order to show that st_data(), which is also part of the julia package, takes extra time on first use because of just-in-time compilation.
On a Windows laptop with an Intel i9-13900H, importing the 100 million data points into a matrix takes just 0.07 seconds for Mata, 0.15 seconds for Julia (the second time), and a substantial 6.73 seconds for Python. In jl, the data copying is performed by C++ code that is multithreaded across columns.
@@
The Julia function is especially slow on first use because that is when @tturbo performs its magic. On later calls, Julia is fast: it takes 0.05 seconds, which works out to about 40 billion floating point operations per second. Mata takes 0.31 seconds and Python, 0.48. To reduce the burden of the first call, if the Julia XQX() function were part of a shared package, it could be accompanied by a script much like the test above. In order to trigger immediate compilation, the Julia package manager would run the script after installing the package. Julia would save the generated machine code for future use.

### v3

@@ Abstract
The julia package integrates the Julia programming language into Stata. Users can transfer data between Stata and Julia, issue Julia commands to analyze and plot, and pass results back to Stata. Julia’s econometric ecosystem is not as mature as Stata’s or R’s or Python’s. But Julia is an excellent environment for developing high-performance numerical applications, which can then be called from many platforms. For example, the boottest program for wild bootstrap–based inference (Roodman et al., 2019) and fwildclusterboot for R (Fischer and Roodman, 2021) can both call the same Julia back end. And the program reghdfejl mimics reghdfe (Correia, 2016) in fitting linear models with high-dimensional fixed effects but calls a Julia package for tenfold acceleration on hard problems. reghdfejl also supports nonlinear fixed-effect models that cannot otherwise be fit in Stata—though preliminarily, as the Julia package for that purpose is immature.
@@ 1 Introduction
In Stata, for example, a call to regress or gsem starts out in the ado programming language. The ado code translates your command into a digital problem definition and passes that to routines written in C/C++ or Stata’s built-in Mata language. In modern Stata, an ado program can even call Python or Java. The called code will probably invoke the LAPACK library, which is written in Fortran, to invert matrices or solve linear systems.
@@
For creators of computationally intensive software, Julia has many strengths. More so than with Stata (but less so than with Python), if you need a tool to perform a fundamental task such as exploiting GPUs, there’s a good chance someone has written and posted it. Julia’s creators set out to solve the “two-language problem,” the inefficiency that arises when programmers have to split a project between high- and low-level languages as they trade off ease of software development and maintenance for speed of execution. (Python and C++ is a common two-language combination, as is R and C++.) Julia resembles Python in syntax, but was built from the start for just-in-time compilation, meaning that it translates Julia functions to machine language when they are first called, and stores that machine code for reuse. A final strength is that, like LISP, Julia can treat code as data. This feature opens the door to meta-programs that can automatically optimize one’s code in certain ways by rewriting it before compilation, or generate new code to take derivatives of complex calculations already implemented.
@@
Julia has disadvantages. While the Julia ecosystem contains the foundation for statistical work—packages to manage data sets, fit standard models, and generate plots—many packages are immature or poorly documented. Standards for storing and returning results are less well developed than in Stata. Much less statistical functionality is available than in Stata or R, or even Python.
@@
Since Julia is a great home for implementing and optimizing numerical software, but not (yet) a great home for the end user, a sweet spot for the new language is back end development. Methods of estimation and inference can be implemented in Julia, and called from Stata, R, and other platforms. The core work can augmented and optimized in one place, rather than separately on each platform.
@@
The julia package links Stata to Julia. It follows in the footsteps of projects to link statistical software to other languages. StataCorp introduced the Stata plugin interface (SPI) for C/C++ in Stata 8.1, in 2003. It added Mata in Stata 9. Fiedler (2012) envisioned running Python inside of Stata. Following the suggestion of a StataCorp employee, Fiedler (2013) partly realized that vision by writing a C++ plugin for Stata. Perhaps inspired by Fiedler, Stata 16 fully integrated Python. Java plugins became possible in Stata 13 and the Java prompt debuted in the Stata results window in version 17. Haghish (2019) introduced rcall, which allows one to issue R commands from Stata; it passes content between the two systems by saving files to disk. Outside of Stata, the JuliaCall (Li, 2019) and JuliaConnectoR (JuliaConnectoR) packages link R to Julia. Another package named JuliaCall does the same for Python. [^1] See juliapy.github.io/PythonCall.jl/stable.
@@
The julia package was most inspired by Fiedler (2013). Its core too is a C++ plugin, which comes precompiled for Windows, Linux, and macOS with Intel or Apple silicon. From the user’s point of view, the package has three components:
@@
The package is designed to hide the complexities of the Stata-Julia linkage, so that getting started requires just two quick steps, installing Julia and installing the julia package for Stata.
Two user-written Stata packages call Julia via julia. The boottest program for wild bootstrap–based inference (Roodman et al., 2019) now accepts a julia option, which causes it to use my Julia package WildBootTests.jl rather than the Mata implementation I originally wrote for boottest. And the new reghdfejl mimics reghdfe and ivreghdfe, programs for fitting linear models with high-dimensional fixed effects (HDFE) (Correia, 2016). reghdfejl calls the Julia package FixedEffectModels.jl, whose lead developer is Matthieu Gomez. reghdfejl also has the ability, in beta, to fit nonlinear HDFE models, such as the logit and negative binomial, via Johannes Boehm’s GLFixedEffectModels.jl. In the last case, using Julia brings new functionality to Stata. In all cases, calling Julia cuts runtime on hard problems, sometimes by an order of magnitude.
Stata users can thus benefit from Julia and julia without using them directly. Still, this paper explains the new julia package. It starts by demonstrating how to run Julia from Stata, in section 2. Section 3 demonstrates applications: it introduces reghdfejl and the Julia back end for boottest; and it demonstrates graphing via Julia. Section 4 gives tips and warns of some limitations. Section 5 more thoroughly describes the package’s Stata commands and Julia functions.
@@ 2.1 Installation and first examples
The preferred way to get Julia is to download the juliaup installation manager. You do that in Windows by installing Julia (not juliaup) from the Windows Store. [^3] The instructions might change. See github.com/JuliaLang/juliaup#installation for the latest. In Linux and macOS, you can instead start Stata and type
@@
This operation may take several minutes. In Windows, many mysterious console windows may appear and disappear as required libraries are downloaded and pre-compiled. In macOS and Linux, output from the installation process may appear in terminal windows. For on first use, jl installs two Julia packages, DataFrames.jl and CategoricalArrays.jl, along with all the packages they depend on. DataFrames.jl (Bouchet-Valat and Kamiński, 2023) is the dominant Julia framework for data sets, i.e., rectangular grids whose columns have names and potentially different data types. CategoricalArrays.jl implements the Julia equivalent of Stata variables with data labels—ones coded, for example, so that 0 means “black” and 1 means “white.” Both packages are hosted on GitHub and are maintained by users, who accept improvements from others.
In general when using Julia, you will experience delays at four points: when packages are installed; when they are first used; when they are first used in a Julia session; and when particular features within them are first used, such as clustering standard errors. A thrust of recent development in the Julia language has been to move the behind-the-scenes processing to earlier stages in order to shorten the lags during regular use.
In the session below, I load Stata’s “auto” data set, extract five variables, run a regression on them, then copy the data to a Julia DataFrame named auto:
@@
jl save is not a Julia command. It is a subcommand of the Stata command jl, which is part of the julia package. If an object named auto already exists in the Julia environment, it is overwritten by the incoming copy of the Stata variables.
Next I type jl by itself, which changes the prompt the same way that typing mata or python or java does. It starts an interactive Julia session, in which I can type multiple Julia commands. I start with displaying the new Julia DataFrame by typing its name:
@@
Along the top are the column names, and below them the types. Int16 means a 16-bit integer, equivalent to Stata int. Float32 corresponds to Stata float. The question marks at the ends of the type names indicate that the types have been modified to allow the special value missing. By default, jl save made all the columns accept missing.
jl save also converted the foreign variable, which is coded as 0/1 in Stata and has value labels “Domestic” and “Foreign,” to a CategoricalVector. While this Julia version of the variable is also stored as a column of integers—1’s and 2’s—it presents as a text variable. [^4] Julia categorical variables are not quite comparable to Stata variables with data labels. Commands involving the Julia variable refer to display values such as “Domestic,” not to numeric codes. The type name “CategoricalVector?” does not fit in the column head, so it is shortened to “Cat…?”.
Next I run the same regression as before, but now in Julia, with the reg() function of the Julia package FixedEffectModels.jl:
  • jl> . jl AddPkg FixedEffectModels  # assure FixedffectModels installed
    jl> . using FixedEffectModels      # load it for use
@@
  • Stata has the # and ## interaction operators. The corresponding symbols in Julia are & and *. The first appears above.
  • All variable names in the command line are prefixed with “:”. That indicates that the names are symbols, not stand-alone Julia objects like the DataFrame auto. The symbols are names of columns in auto.
  • Because foreign is stored as a CategoricalVector, reg() automatically treats it as a factor variable, generating a pair of dummies to be interacted with mpg. There is no need for a Julia equivalent of Stata’s “i.” prefix. However, in Julia, treating non-CategoricalVectors as factor variables requires a more awkward syntax. One must pass reg() an argument with a dictionary mapping the variables to coding types, such as with contrasts=Dict(:headroom=>DummyCoding()).
The command line contains deeper complexity. In Julia, as in R and Python, every regression command is a statement in a full-fledged programming language. This tends to complicate the syntax for even basic operations. It is hard to implement an R or Python line as clean as regress y x. In the Julia command above, term(:price), term(:headroom), …, etc., create instances of the object type Term to represent variables within a statistical model. The Term structure is defined in the StatsModels.jl package, which FixedEffectModels.jl loads and makes available to the user. The content of each Term is just a symbol, such as :price. What give Term’s life are the operations that StatsModels.jl defines on them. The + operator is repurposed (“overloaded”) to form lists (tuples) of Term’s. The $\sim$ operator binds together two Term tuples to make a FormulaTerm, a structure whose distinctive feature is having designated left and right sides, just as in a linear regression model. In the example, an entire FormulaTerm is passed as the second argument to reg(). In this way, the primitives of a flexible programming language are composed to communicate a statistical model.
This standard for representing regression models is not intrinsic to Julia. It has emerged from the work of users. And it is transparent. In Stata, one could type jl: dump(term(:c) $\sim$ term(:a) + term(:b)) to reveal the contents of a FormulaTerm. The software that builds and applies these structures is on GitHub, all written in Julia. The StatsModels.jl formula language is also extensible [^5] See juliastats.org/StatsModels.jl/stable/internals/#extending.. As we will see, FixedEffectModels.jl extends the formula language in order to mark variables whose fixed effects are to be absorbed.
Yet a powerful feature of Julia makes it practical to ignore most of this complexity. The regression command above can be simplified with a macro. Where macros in Stata are strings, in Julia they are code for manipulating code: meta-programs. StatsModels.jl defines a macro called @formula, which lets one write models more simply. Thus:
@@
Before this new command is compiled and executed, @formula turns it into something closer to the earlier regression command. [^6] Johannes Boehm’s experimental Douglass.jl package includes macros to translate Stata-like commands into Julia ones. An example: d"bysort :Species : egen :z = sum(:SepalLength) if :SepalWidth .> 3.0".
Notice one other change to the regression command: it now starts with m = in order to store the fitted model in the Julia variable m. This lets me extract return values and pass them back to Stata:
@@ 2.2 Built for speed
To partially convey why Julia is a good environment for scientific computing, I present another example. Suppose we have a data set with 10 million rows and 10 columns named x1-x10. Given a $10\times 10$ matrix ${\bf Q}$, we want to compute the norm of each row according to the quadratic form defined by ${\bf Q}$ and store the results in a new column. That is, for each row $x_{i}$, now viewed as a column vector, we want $x_{i}^{\prime}{\bf Q}x_{i}$. It is natural to implement this computation with triply nested for loops. That’s a good way to do it in Julia (inside Stata):
@@
We saw in the previous example that one can start an interactive Julia session inside Stata by typing jl (and end it with exit()). But the interactive mode is not accessible to do files. To exemplify Julia code in a do file, the above snippet is therefore written effectively as a single line. It is typographically, not syntactically, split across multiple lines with Stata’s continuation token, ///. When putting many Julia commands on what is syntactically one line, most commands not immediately followed by end need to terminate with a semicolon, as illustrated above.
The code for this new Julia function, XQX(), is mostly straightforward. It initializes the return value to a column of 0’s, performs the calculation through nested loops, and returns the result. One detail is unusual: another macro is invoked. @tturbo is part of Chris Elrod’s LoopVectorization.jl package. @tturbo analyzes the code that follows it and substantially—but invisibly—rewrites it for speed. It may unroll a loop to reduce the number of jumps back to the top. The macro, or the compiler, may spot the opportunity to move X[i,j] * out of the inner loop since it does not depend on the loop’s index, k. The macro may even reorder the nesting of the loops with an eye toward the fact that in Julia, matrices are stored column by column, and on-chip memory caching makes it is faster to access adjacent memory locations in sequence. It may tailor the code so that the resulting machine language exploits the ability of modern chips to vectorize operations within each CPU core, in 256- or 512-bit registers—what is called single instruction, multiple data (SIMD) execution. And @tturbo will exploit multithreading, the subdivision of work across cores; that is what the first “t” in tturbo stands for.
Triply nested implementations in Mata and Python are much slower, I assume because the calculations in the innermost loop have to be interpreted a billion times. In constrast, Julia translates all the code down to machine language just once. The best alternatives I have found in the other two languages are essentially the same:
@@
Both delegate the computation to built-in matrix operations, which are fast because they are implemented in compiled C or Fortran. However, both contain a subtle inefficiency. They create the large, temporary matrices ${\bf XQ}$ and ${\bf XQ\otimes\bf X}$, before summing their rows and dispensing with them. Since in our example ${\bf X}$ is $10^{7}\times 10$, and since double-precision numbers take 8 bytes, the temporary matrices require some 800 megabytes each. Allocating and deallocating memory itself takes time. Expanding the memory footprint of an algorithm can also reduce speed by exceeding the capacity of a CPU’s memory caches, forcing the CPU to idle while data are written to and read from the computer’s main memory. In contrast, the triply-looped Julia code demands essentially no temporary storage.
This inefficiency of Mata and Python can be seen as an artifact of the two-language problem Julia was designed to solve. The best one can do in those languages is rely on a set of canned routines, such as for matrix multiplication, that are written in another language (C++ or Fortran), routines that, while fast on their own terms, are nearly impossible to revise, yet not optimized for the purpose at hand.
To create a test bed for these functions, I construct ${\bf X}$ as a Stata data set and ${\bf Q}$ as a Stata matrix, then copy them into Mata, Python, and Julia. The set rmsg on command generates timing reports:
@@
Under a Julia convention, the semicolons at the end of the jl lines suppress printing of the results of the assignments. The Julia data copying command is run twice in order to show that st_data(), which is also part of the julia package, takes extra time on first use because that is when it is compiled down to machine code.
On a Windows laptop with an Intel i9-13900H processor, importing the 100 million data points into a matrix takes just 0.07 seconds for Mata, 0.15 seconds for Julia (the second time), and a substantial 6.73 seconds for Python. In jl, the data copying is performed by C++ code that is multithreaded across columns.
@@
The Julia function is especially slow on first use because that is when @tturbo performs its magic. On later calls, Julia is fast: it takes 0.05 seconds, which works out to about 40 billion floating point operations per second. Mata takes 0.31 seconds and Python, 0.48. To reduce the burden of the first call, if the Julia XQX() function were part of a shared package, it could be accompanied by a script much like the test above. The Julia package manager would run the script after installing the package in order to trigger immediate compilation. Julia would save the generated machine code for future use.